OpenStudy (anonymous):

What is equal to 2/4

OpenStudy (lgbasallote):

\[\frac 24 = \frac 12 = \frac 36 = \frac 48 = \frac {5}{10} = \cdots\] etc,

OpenStudy (lgbasallote):

you just keep multiplying/dividing the numerator and denominator by the same number

OpenStudy (saifoo.khan):

Sometimes 0.5 sometimes 1/2

OpenStudy (unklerhaukus):

50%